OpenStudy (anonymous):

What is the sum of all values of N, 1≤N≤999, such that 1+2+3+…+N is a perfect square?

OpenStudy (anonymous):

The sum n n numbers from 1 to n is given by: \[\frac{n(n+1)}{2}\] Are you asking for a value of N such that the sum is a perfect square, or for the sum of numbers from 1 to 999?

OpenStudy (chihiroasleaf):

if n = 1 , n = 8

OpenStudy (anonymous):

there are many numbers for N i am asking their sums

OpenStudy (anonymous):

@Traxter gave you the sums of natural numbers

OpenStudy (anonymous):

It might be helpful to note that the sum of n odd numbers is a perfect square: \[1+3+5+...+(2n-1)=n^2\] So we want the sum of all the odd numbers betwee 1 and 999. So take (2n-1)=999, so that 2n=1000 and n=500. So The sum of all the odd numbers between 1 and 999 is \(500^2\)=250,000.

OpenStudy (anonymous):

I think that's what you're looking for @irudayadhason.

OpenStudy (anonymous):

the answer should be an integer between 0-999

OpenStudy (anonymous):

The question really doesn't make much sense, do you think you could rephrase it please?

OpenStudy (anonymous):

no there are no mistakes in this question

OpenStudy (anonymous):

i think he is looking for values of N but all the answers added up @irudayadhason

OpenStudy (anonymous):

so we want to solve\[1+2+3+...+n=m^2\]\[n(n+1)=2m^2\]

OpenStudy (anonymous):

no i am not looking for the values of N i am just asking the sum of all values of N but it should be between 0-999

OpenStudy (asnaseer):

Maybe this will help explain it to others better: lets say the only solutions to this were n=1 and n=8, then your answer would be 1 + 8 = 9. so, from all the valid solutions to this (\(n_1,n_2,n_3,...\)), the answer we want is:\[n_1+n_2+n_3+...\]

OpenStudy (anonymous):

sorry 9 is a wrong answer try to use this key technique "Pell's Equation."

OpenStudy (asnaseer):

I didn't say 9 way the answer - I was using it as an example to explain the question. :)

OpenStudy (asnaseer):

*was

OpenStudy (asnaseer):

the aim of this site is not "just to hand out answers". it is to teach and learn.

OpenStudy (anonymous):

\[n(n+1)=2m^2\]\[4n(n+1)=8m^2\]\[(2n+1)^2=8m^2+1\]letting\[x=2n+1\]\[y=2m\]will gives\[x^2-2y^2=1\]

OpenStudy (anonymous):

you mean the answer is ...

OpenStudy (anonymous):

Traxter please try to solve

OpenStudy (anonymous):

later equation solved by Pell see here http://en.wikipedia.org/wiki/Square_triangular_number

OpenStudy (anonymous):

answer will be\[x=P_{2k}+P_{2k-1}\]\[y=P_{2k}\]\(P_k\) is kth Pell number see here http://en.wikipedia.org/wiki/Pell_number

OpenStudy (anonymous):

i need an answer between 0-999